# Neutral Barium in Solid Neon: Optical Spectroscopy and First Excited State Lifetime

### v1

@@ I Introduction
In recent years, spectroscopy of atoms and molecules embedded into noble gas cryogenic crystals has developed significantly, as it plays a key role for high precision measurements within the matrix isolation technique (MIT). MIT enables the confinement of highly reactive species (dopants) within an inert, low‑temperature host DUBEY2018317, to investigate their characteristics and behavior while interrogating them at a fixed point in space and simultaneously achieving a high density of dopants in the crystal Weinstein2005.
@@
Other techniques to investigate atoms or molecules include slowing a production beam Patterson2007 or trapping them through Magneto Optical Traps (MOTs) Raab1987. These methods enable in-gas investigation of the species, at the cost of the number of molecules. However, for fundamental physics research, having a large number of atoms or molecules enhances the sensitivity of the measurements Safronova2018. Therefore, in this study, we discuss the production of an inert cryogenic environment with a high density of dopants and investigate them through spectroscopic measurements.
@@
The investigation method used is Laser Induced Fluorescence (LIF) as it allows for high signal to noise ratio and good sensitivity for atomic or molecular spectroscopy Kinsey1977.
@@
Neutral barium (Ba) has a favorable electronic structure to be observed in noble gas matrices Callear1973, Trenary1973, and has been recently characterized in matrices such as Ar, Kr and Xe using LIF spectroscopy, revealing multiple occupancy sites and Jahn-Teller splitting McCaffrey, McCaffrey_BaSr_inRGmatrices. Further studies in solid helium have also provide new insights on how barium, under very low perturbation conditions, shows a very long lifetime of atomic metastable states of the orders of seconds Weis_BaHelium.
@@
Figure 1: Top view sketch of the experimental setup including production of a gas phase barium beam, crystal growth, excitation, and fluorescence detection: A) Barium production and doped neon crystal growth setup: two lasers are simultaneously used. An Nd:YAG laser (light blue arrows in the figure) is used for the ablation of the metallic barium target. An Avesta laser (red arrows in the figure) at 12636 cm-1 (791.13 nm) is used to probe the presence of barium atoms, ensuring efficient production both inside and outside the cell. B) Excitation and fluorescence detection scheme: the gas composed of Ba and neon is sprayed onto the CaF2 substrate (orange rectangle) at 6.8 K, which is permanently tilted to minimize the collection of excitation light in the fluorescence detection system. Purple arrows indicate the excitation laser. A custom built lens assembly allows us to fix the first collection lens at about 30 mm from the crystal growth spot, enabling efficient collection of the emitted signal which is then focused on either a spectrometer or a photodiode, depending on the measurement.
@@
In this paper, we extend the approach conducted in previously cited works, proving the first successful growth of good quality Ba doped neon crystals at temperatures ranging from 6.8 to 9 K. Among the hosts previously studied, neon is a promising matrix for MIT, as it can offer good optical transparency and reduced perturbation of embedded species, as proven in previous studies of atoms embedded in cryogenic neon crystals 1Yb_Neon_Lambo, dargyte2021optical_Rb_Neon_Spin, braggio2022spectroscopy_RbNEon. Therefore, this approach enables direct comparison of Ba properties in our neon crystal lattice with respect to the previously studied ones.
@@
In this work, we investigate neutral barium atoms embedded in a neon matrix, where neon serves both as a buffer gas in our atomic source and as a matrix deposition gas for crystal growth. This systems enables LIF spectroscopic measurements of barium atoms in the crystal, using single and double laser excitation in either continuous or pulsed mode. These methods enable selective excitation of barium energy levels. We also measure, for the first time, the lifetime of the metastable state 5d6s 3D1 of barium in a neon matrix.
@@
This approach creates not only a new base for spectroscopic analysis of barium in inert cryogenic matrices, but also an essential experimental framework for embedding molecules, such as barium monofluoride (BaF), in cryogenic crystals. BaF is a molecule with great characteristics for low energy, high precision measurements such as the electron electric dipole moment (eEDM) through the MIT.
@@
BaF production will include deposition of some neutral barium presence in the matrix, unless it is filtered out yau2024specularreflectionpolarmolecules, corriveau2024matrixisolatedbariummonofluoride at the cost of a loss in the number of molecules. Therefore, understanding the spectral signatures and dynamics of Ba will be essential, as it will provide a crucial spectroscopic baseline for future BaF matrix isolation measurements in a cryogenic inert environment such as neon Li_2023, corriveau2024matrixisolatedbariummonofluoride or parahydrogen Messineo_pH, Borghesani_2023, as recently proposed by our group within the DOCET experiment.
@@ II.1 Barium doped neon crystals production
[…]
@@
To probe the production of neutral barium atoms in the gas phase, before deposition, two absorption measurements are performed simultaneously, as already done in previous studies for BaF mooij2024novel, inside and outside of the cell. An Avesta Ti:Sa laser is used, with a linewidth of approximately 2 GHz, tunable over a 14 300 - 11100 cm-1 range (700 - 900 nm). The laser is tuned to 12 636 cm-1 (791.13 nm), probing the transition from the ground state 6s2 $^{1}\text{S}_{0}$ to 6s6p $3\text{P}_{1}$. The barium energy levels used in this work are shown in Figure 2. An absorption measurement probes the production directly inside the cell, where the laser is sent through a window at the back of the cell and exits from the extraction hole towards a Thorlabs DET36A/M silicon photodiode. The second method consist in a differential measurement with a Thorlabs PDB210A/M. To do so, a laser beam is equally split, one part is used as reference signal, while the other passes 5 mm from the extraction hole, in a dual pass configuration, aligned through two diaphragms at the side of the extraction hole. Noise is highly suppressed in this configuration, allowing for the detection of bunches of atoms coming out of the source every ablation pulse.
@@ II.2 Optical setup for fluorescence collection
Light collection occurs almost orthogonally from the excitation beam, to reduce the unwanted excitation signal and optimize the fluorescence detection.
@@ III.1 High energy excitation of barium in neon matrix
The work of Lebedev et al. Weis_BaHelium demonstrate that a third harmonic of an Nd:YAG laser, with pulses at 28 195 cm-1 (355 nm) can excite a wide range of barium states by both radiative or non-radiative transitions in a solid helium matrix. Among these, long lived metastable state of barium, such as 5d6s 1D2 and the 5d6s 3Dj ($j=1,2,3$) manifold, act as an intermediate reservoir, and when a second pulse excites them, they can reach excited state at energies over 30 000 cm-1 with respect to the energy of the ground state. This method leads to barium fluorescence emission through a radiative cascade, as from the highest excited state the atom can populate a metastable state before decaying into the ground state.
We adapted this approach to detect barium fluorescence in a neon crystal (vis-NIR), enabling the first comparison up to date with the work of Lebedev et al. Weis_BaHelium in solid helium.
@@
Table 1: Barium emission transitions in a neon crystal at 6.8 K were observed following excitation with the third harmonic (355 nm) of a Nd:YAG laser. The spectral shifts are referenced against the corresponding free atomic barium lines, listed in the lower portion of the table. Transition H is an electric dipole forbidden line in the gas phase NIST_ASD. Symbol $\ast$ = multiple transitions involved.
@@
We identified a total of eight fluorescence peaks. Comparing our data with the spectroscopic results of Lebedev et al. on barium in solid helium Weis_BaHelium, four fluorescence lines match those previously reported.
In general, the relatively narrow wavelength shifts with respect to the free barium atomic lines and the relatively small full width at half maximum (FWHM) allow us to reliably identify five electronic state transitions (C, E, F, G, H, see Figure 3), as shown in Table 1. Among these, the largest observed shift is for peak C (93-93 cm-1), while the smallest is for peak F (9-9 cm-1). The magnitude of these shifts is smaller than those reported for barium isolated in rare gas matrices McCaffrey.
@@
The FWHM values can be directly compared with those obtained by Lebedev et al. Weis_BaHelium of barium in solid helium at 1.5 K. Peak C, corresponding to the 6s6p $^{1}\text{P}_{1}$ to 6s2 $^{1}\text{S}_{0}$ transition, has a width of $\Delta\nu=156\ \text{cm}^{-1}$, broader than the helium value ($\delta\nu=30\ \text{cm}^{-1}$). However, peak E, F and G show the same width with respect to the same transitions observed in solid helium.
@@ III.2 Excitation with single and double laser
In this section, our objective is to investigate the doped crystal using continuous-wave lasers, following the work of Li et al. Li_2023. To this end, we irradiated 1 mm-wide spots on the crystal surface with Ti:Sa lasers delivering up to 100 mW, and measured excitation spectra by tuning the laser over its operational range of 14 300–11 100 cm-1 (700–870 nm).
@@
These excitation wavelengths are far detuned from direct transitions and likely populate the 6s6p 3Pj manifold through matrix-induced relaxation, as shown in Figure 4. Those states can decay to the ground state through radiative cascade involving the 5d6s 3Di (i = 1,2,3) and 1D2 states. Direct fluorescence from the 3Pj states is not detected, as it spectrally overlaps with the excitation light. Our working hypothesis is that the excitation laser directly addresses the 6s6p 3P2 level at 13 514.7 cm-1 (740.0 nm), a strongly suppressed transition in the free atom. However, in the matrix environment, the local perturbation breaks the atomic symmetries that govern the selection rules in the gas phase moroshkin2008atomic. As a result, transitions that are forbidden in free barium become weakly allowed, allowing the population transfer to the 6s6p 3P2 state and the subsequent relaxation into the 5d6s 3Di manifold or 5d6s 1D2 state.
As stated in precedent spectroscopy studies of Ba atoms in RG gas matrices, the line broadening of excitation is larger with respect to the fluorescence emission. In this case, the effect is likely enhanced by the indirect excitation discussed above.
@@
Interestingly, we also detected green fluorescence around 18000 cm-1 corresponding to the 6s6p 1P1 state, as shown in Figure 7 (a).
@@
Interestingly the 6s6p 3P1 state exhibits two dominant fluorescence bands centered at 18 622 (557 nm) and 17 967 (538 nm) cm-1. Similar features have been extensively studied in Ar, Kr, and Xe matrices McCaffrey, where they are attributed to the influence of lattice site occupancy on barium emission properties.
Figure 8: Energy level diagram (not to scale) of the proposed double laser pumping scheme for barium in a neon matrix. The first laser (red arrow) at 13 369 cm-1 (749 nm) indirectly populates the 5d6s 3D2 state. From this level, a second laser (blue arrow) tuned at 12 849 cm-1 (780 nm) excites the atom to the higher 5d6p 3F2 state. According to the cascade dynamics reported in previous works Weis_BaHelium, this state decays to the 5d6s 1D2 level with emission at 10 670 cm-1 (937 nm), followed by a second fluorescence decay to the ground state 6s2 1S0 at 11 394 cm-1 (877 nm). The black arrows represent spontaneous emission pathways, while red arrows denote laser driven excitations. Dashed lines indicate indirect population routes from the 6s6p 3Pj manifold.
We observe a high intensity fluorescence emission when the excitation laser is tuned to approximately 14 180 cm-1 (705 nm). However, it does not correspond to any neutral atomic barium state, and no direct excitation pathway from the ground or low lying metastable states to higher lying electronic states has been identified at this wavelength. This discrepancy suggests that the observed fluorescence may arise from matrix induced perturbations, site selective effects, or indirect population mechanisms through long lived intermediate states. Further investigation is beyond the scope of this work.
To further test the feasibility of the double–laser excitation, we designed a pumping scheme path that populates 5d6p 3F2 at 22 065 cm-1 (453 nm) only if we strongly populate the manifold 6s6d $^{3}\text{D}_{j}$, as shown in Figure 8. Although the excitation bands are $\approx$1000 cm-1 wide, our measurements demonstrate that this scheme works exclusively via indirect population of the first manifold, specifically through the 5d6s 3D2 state. A second laser, set at 12 849 cm-1, can then couple this state and populate the 5d6p 3F2 state laying at 22 064.622\,064.6 cm-1 (453 nm). This scheme relays on the evidence shown earlier that excitation near 13 369 cm-1 (748 nm) populates the 5d6s 3D1 and hypothetically the other states of the same 5d6s 3Dj manifold.
@@
To further test the scheme, if we tune the first laser to 13 504 cm-1 (740 nm), a known population maximum of the 5d6s 1D2 state, no excitation pathway to the 5d6p 3F2 level is observed, as shown in Figure 9 (b), probing our selective pumping scheme functioning.

### v2

@@ I Introduction
In recent years, spectroscopy of atoms and molecules embedded into noble gas cryogenic crystals has developed significantly, as it plays a key role for high precision measurements within the matrix isolation technique (MIT). MIT enables the confinement of highly reactive species (dopants) within an inert, low‑temperature host Dubey et al. [2018], to investigate their characteristics and behavior while interrogating them at a fixed point in space and simultaneously achieving a high density of dopants in the crystal Upadhyay et al. [2020].
@@
Other techniques to investigate atoms or molecules include slowing a production beam Patterson and Doyle [2007] or trapping them through Magneto Optical Traps (MOTs) Raab et al. [1987]. These methods enable in-gas investigation of the species, at the cost of the number of molecules. However, for fundamental physics research, having a large number of atoms or molecules enhances the sensitivity of the measurements Safronova et al. [2018]. Therefore, in this study, we discuss the production of an inert cryogenic environment with a high density of barium (Ba) atoms as dopants to investigate them through spectroscopic measurements as a testbench for future studies of more complex molecules such as barium monofluoride (BaF).
@@
The investigation method used is Laser Induced Fluorescence (LIF) as it allows for high signal to noise ratio and good sensitivity for atomic or molecular spectroscopy Kinsey [1977].
@@
Neutral barium has a favorable electronic structure to be observed in noble gas matrices Callear and Metcalfe [1973], Trenary and Shriver [1973], and has been recently characterized in matrices such as Ar, Kr and Xe using LIF spectroscopy, revealing multiple occupancy sites and Jahn-Teller splitting Davis and McCaffrey [2018, 2016]. Further studies in solid helium have also provide new insights on how barium, under very low perturbation conditions, shows a very long lifetime of atomic metastable states of the orders of seconds Lebedev et al. [2011].
@@
Figure 1: Top view sketch of the experimental setup including production of a gas phase barium beam, crystal growth, excitation, and fluorescence detection: A) Ba doped neon crystal growth setup: two lasers are simultaneously used. An Nd:YAG laser (light blue arrows in the figure) is used for the ablation of the metallic barium target, which gets extracted from the cell through a hole in front of a CaF2 substrate. An Avesta laser (red arrows in the figure) at 12636 cm-1 (791.13 nm) is used to ensure efficient production both inside and outside the cell during ablation. B) Excitation and fluorescence detection scheme: the gas composed of Ba and neon is sprayed onto the CaF2 substrate (orange rectangle) at 6.8 K, which is permanently tilted to minimize the collection of excitation light (purple arrows) in the fluorescence detection system. A custom built lens assembly allows us to fix the first collection lens at about 30 mm from the crystal growth spot, enabling efficient collection of the emitted signal which is then focused on either a spectrometer or a photodiode.
@@
In this paper, we extend the approach conducted in previously cited works, proving the first successful growth of good quality Ba doped neon crystals at temperatures ranging from 6.8 to 9 K. Among the hosts previously studied, neon is a promising matrix for MIT, as it can offer good optical transparency and reduced perturbation of embedded species, as proven in previous studies of atoms embedded in cryogenic neon crystals Lambo et al. [2012], Dargyte et al. [2021], Braggio et al. [2022]. Therefore, this approach enables direct comparison of Ba properties in our neon crystal lattice with respect to the previously studied ones.
@@
In this work, neon serves both as a buffer gas in our atomic source and as a matrix deposition gas for crystal growth. Our setup permits LIF spectroscopic measurements of barium atoms in the crystal, using single and double laser excitation in either continuous or pulsed mode. These methods enable selective excitation of barium energy levels. We also measure, for the first time, the lifetime of the metastable state 5d6s 3D1 of barium in a neon matrix.
@@
The study of barium in cryogenic inert matrices serves as an initial spectroscopic approach towards molecules, such as BaF, with great characteristics for low energy, high precision measurements such as the electron electric dipole moment (eEDM) through the MIT.
@@
BaF production will include deposition of some neutral barium presence in the matrix, unless it is filtered out Yau et al. [2024], Corriveau et al. [2024] at the cost of a loss in the number of molecules. Therefore, understanding the spectral signatures and dynamics of Ba will be essential, as it will provide a crucial spectroscopic baseline for future BaF matrix isolation measurements in a cryogenic inert environment such as neon Li et al. [2023], Corriveau et al. [2024] or parahydrogen Antonini et al. [2025], Borghesani et al. [2023], as recently proposed by our group within the DOCET experiment.
@@ II.1 Barium doped neon crystals production
Neutral barium atoms are produced in a custom designed stainless steel cryogenic ablation cell, place in a vacuum chamber maintained below $1\times 10^{-7}$ mBar as shown in Figure 1 A), exploiting the buffer gas cooling technique Doyle et al. [1995], Bethlem et al. [2002]. Laser ablation, within this method, allows for the production of large numbers of atoms at very low temperature, as collisions with the cold buffer gas cool them down. In our setup, neon is used as a buffer gas. The cell is cooled with a pulse tube refrigerator Leybold-Heraeus RG210 and kept at approximately 20 K, to avoid neon condensation while still allowing efficient cooling of the produced atoms.
@@
To probe the production of neutral barium atoms in the gas phase, before deposition, two absorption measurements are performed simultaneously, as already done in previous studies for BaF Mooij et al. [2024], inside and outside of the cell. An Avesta Ti:Sa laser is used, with a linewidth of approximately 2 GHz, tunable over a 14 300 - 11100 cm-1 range (700 - 900 nm). The laser is tuned to 12 636 cm-1 (791.13 nm), probing the transition from the ground state 6s2 $^{1}\text{S}_{0}$ to 6s6p $3\text{P}_{1}$. The barium energy levels used in this work are shown in Figure 2. An absorption measurement probes the production directly inside the cell, where the laser is sent through a window at the back of the cell and exits from the extraction hole towards a Thorlabs DET36A/M silicon photodiode. The second method consist in a differential measurement with a Thorlabs PDB210A/M. To do so, a laser beam is equally split, one part is used as reference signal, while the other passes 5 mm from the extraction hole, in a dual pass configuration, aligned through two diaphragms at the side of the extraction hole. Noise is highly suppressed in this configuration, allowing for the detection of bunches of atoms coming out of the source every ablation pulse.
@@ II.2 Optical setup for fluorescence collection
Light collection occurs almost orthogonally from the excitation beam, as illustrated in Figure 1 in purple arrows, to reduce the unwanted excitation signal and optimize the fluorescence detection.
@@ III.1 High energy excitation of barium in neon matrix
The work of Lebedev et al. Lebedev et al. [2011] demonstrate that a third harmonic of an Nd:YAG laser, with pulses at 28 195 cm-1 (355 nm) can excite a wide range of barium states by both radiative or non-radiative transitions in a solid helium matrix. Among these, long lived metastable state of barium, such as 5d6s 1D2 and the 5d6s 3Dj ($j=1,2,3$) manifold, act as an intermediate reservoir, and when a second pulse excites them, they can reach excited state at energies over 30 000 cm-1 with respect to the energy of the ground state. This method leads to barium fluorescence emission through a radiative cascade, as from the highest excited state the atom can populate a metastable state before decaying into the ground state.
We adapted this approach to detect barium fluorescence in a neon crystal (vis-NIR), enabling the first comparison up to date with the work of Lebedev et al. Lebedev et al. [2011] in solid helium.
@@
Table 1: Barium emission transitions in a neon crystal at 6.8 K were observed following excitation with the third harmonic (355 nm) of a Nd:YAG laser. The spectral shifts are referenced against the corresponding free atomic barium lines, listed in the lower portion of the table. Transition H is an electric dipole forbidden line in the gas phase Kramida et al. [2024]. Symbol $\ast$ = multiple transitions involved.
@@
We identified a total of eight fluorescence peaks. Comparing our data with the spectroscopic results of Lebedev et al. on barium in solid helium Lebedev et al. [2011], four fluorescence lines match those previously reported.
In general, the relatively narrow wavelength shifts with respect to the free barium atomic lines and the relatively small full width at half maximum (FWHM) allow us to reliably identify five electronic state transitions (C, E, F, G, H, see Figure 3), as shown in Table 1. Among these, the largest observed shift is for peak C (93-93 cm-1), while the smallest is for peak F (9-9 cm-1). The magnitude of these shifts is smaller than those reported for barium isolated in rare gas matrices Davis and McCaffrey [2018].
@@
The FWHM values can be directly compared with those obtained by Lebedev et al. Lebedev et al. [2011] of barium in solid helium at 1.5 K. Peak C, corresponding to the 6s6p $^{1}\text{P}_{1}$ to 6s2 $^{1}\text{S}_{0}$ transition, has a width of $\Delta\nu=156\ \text{cm}^{-1}$, broader than the helium value ($\delta\nu=30\ \text{cm}^{-1}$). However, peak E, F and G show the same width with respect to the same transitions observed in solid helium.
@@ III.2 Excitation with single and double laser
In this section, our objective is to investigate the doped crystal using continuous-wave lasers, following the work of Li et al. Li et al. [2023]. To this end, we irradiated 1 mm-wide spots on the crystal surface with Ti:Sa lasers delivering up to 100 mW, and measured excitation spectra by tuning the laser over its operational range of 14 300–11 100 cm-1 (700–870 nm).
@@
These excitation wavelengths are far detuned from direct transitions and likely populate the 6s6p 3Pj manifold through matrix-induced relaxation, as shown in Figure 4. Those states can decay to the ground state through radiative cascade involving the 5d6s 3Di (i = 1,2,3) and 1D2 states. Direct fluorescence from the 3Pj states is not detected, as it might be too close to the excitation light which was properly filtered out. Stoke shifts are usually big enough to directly detect emission light, but in neon such transitions have not been detected in the spectral region of collection. In the matrix environment, the local perturbation breaks the atomic symmetries that govern the selection rules in the gas phase Moroshkin et al. [2008]. As a result, transitions that are forbidden for free barium become weakly allowed. In our case, the excitation laser seems to directly address the 6s6p 3P2 level at 13 514.7 cm-1 (740.0 nm), a strongly suppressed transition in the free atom, allowing the population transfer to the 6s6p 3P2 state and the subsequent relaxation into the 5d6s 3Di manifold or 5d6s 1D2 state.
As stated in precedent spectroscopy studies of Ba atoms in RG gas matrices Davis and McCaffrey [2018], the line broadening of excitation is larger with respect to the fluorescence emission. In this case, the effect is likely enhanced by the indirect excitation discussed above.
@@
Green fluorescence around 18000 cm-1 corresponding to the 6s6p 1P1 state has been detected, as shown in Figure 7 (a).
@@
The 6s6p 3P1 state exhibits two dominant fluorescence bands centered at 18 622 (557 nm) and 17 967 (538 nm) cm-1. Similar features have been extensively studied in Ar, Kr, and Xe matrices Davis and McCaffrey [2018], where they are attributed to the influence of lattice site occupancy on barium emission properties. Further investigation of the process of site selection of Barium in a neon crystal is out of the scope of our project and has already been discussed for BaF embedded in a neon matrix Lambo et al. [2023].
Figure 8: Energy level diagram (not to scale) of the proposed double laser pumping scheme for barium in a neon matrix. The first laser (red arrow) at 13 369 cm-1 (749 nm) indirectly populates the 5d6s 3D2 state. From this level, a second laser (blue arrow) tuned at 12 849 cm-1 (780 nm) excites the atom to the higher 5d6p 3F2 state. According to the cascade dynamics reported in previous works Lebedev et al. [2011], this state decays to the 5d6s 1D2 level with emission at 10 670 cm-1 (937 nm), followed by a second fluorescence decay to the ground state 6s2 1S0 at 11 394 cm-1 (877 nm). The black arrows represent spontaneous emission pathways, while red arrows denote laser driven excitations. Dashed lines indicate indirect population routes from the 6s6p 3Pj manifold.
We observe a high intensity fluorescence emission when the excitation laser is tuned to approximately 14 180 cm-1 (705 nm) which is probably populating the 5d6s 1D2 state as shown in Figure 5, as well as two dominant emission when a second laser excites the sample at around 12 270 cm-1 (815 nm) and 12 060 cm-1 (830 nm). We note that the three excitations are close to BaF energy levels reported in previous works Li et al. [2023]. The observed emission is originating from different population pathways involving long living intermediate states, as shown in the following sections.
To further test the feasibility of the double–laser excitation, we designed a pumping scheme path that populates 5d6p 3F2 at 22 065 cm-1 (453 nm) only if we strongly populate the manifold 6s6d $^{3}\text{D}_{j}$, as shown in Figure 8. Although the excitation bands are $\approx$1000 cm-1 wide, our measurements demonstrate that this scheme works exclusively via indirect population of the first manifold, specifically through the 5d6s 3D2 state. A second laser, set at 12 849 cm-1, can then couple this state and populate the 5d6p 3F2 state laying at 22 06522\,065 cm-1 (453 nm). This scheme relays on the evidence shown earlier that excitation near 13 369 cm-1 (748 nm) populates the 5d6s 3D1 and hypothetically the other states of the same 5d6s 3Dj manifold.
@@
To further test the scheme, if we tune the first laser to 13 504 cm-1 (740 nm), a known population maximum of the 5d6s 1D2 state, no excitation pathway to the 5d6p 3F2 level is observed, as shown in Figure 9 (b), probing our selective pumping scheme functioning. Moreover, both lasers have been applied to the crystal at higher intensities, individually and simultaneously, as saturation is usually never reached in a matrix isolation environment and additional atom emissions might arise. However, the obtained spectra exhibit the same characteristic as those reported here.